# Perception-Distortion Balanced Super-Resolution: A Multi-Objective Optimization Perspective

### v1

@@ IV-C Ablation Studies
To validate the effectiveness of the proposed hybrid EA-Adam optimization and network fusion methods, we conduct experiments to compare our hybrid EA-Adam algorithm with the Adam algorithm, and compare our network fusion method with the learnable weight method. Furthermore, different loss functions are applied to optimize the weight regression network to show the influence of loss selection. We also study the weights of obtained models with different perception-distortion preferences. All ablation results are evaluated on the Urban100 dataset.
@@
Model
Model1
Model2
Model3
Model4
Fused model
Weights
$[0.25,0.75]$
$[0.5,0.5]$
$[0.75,0.25]$
$[1,0]$
PSNR
Adam
24.36
23.82
23.96
24.22
24.41
EA-Adam
25.78 25.75 24.95 24.73 24.83
AdamW
25.48
24.67
23.41
24.18
24.37
EA-AdamW
25.95 25.81 25.21 24.82 24.91
LPIPS
Adam
0.3061
0.2899
0.2665
0.1446
0.1437
EA-Adam
0.2120 0.2055 0.1536 0.1384 0.1354
AdamW
0.2570
0.2925
0.2306
0.1477
0.1444
EA-AdamW
0.2265 0.2201 0.1519 0.1404 0.1368
@@
Table IV gives the quantitative comparison. Without loss of generality, we conduct experiments with the SRResNet backbone and evaluate on the Urban100 dataset. PSNR and LPIPS metrics are used to measure the fidelity and perceptual quality, respectively. One can see that our hybrid EA-Adam/AdamW optimizers outperform Adam/AdamW by a large margin in most cases. Both the fidelity and perceptual quality metrics are improved, validating that the proposed hybrid EA-Adam strategy can push the PF closer to the optimal point of the metric space. For most models, EA-Adam improves PSNR by more than 0.4dB over Adam, and improves LPIPS by more than $10\%$. This might be caused by the gradient vanishing problem in the gradient-based optimization process so that the models can be trapped into a local minimum. In contrast, with our hybrid EA-Adam optimizer, effective interaction between different models can be introduced by the crossover operator in EA, so that the models can be stably optimized with better convergence.
On the other hand, each of the four models may perform well in one aspect, yet sacrifices the other aspect as a trade-off. The fused model by our network fusion method achieves a better balance between fidelity and perceptual quality. The LPIPS of the fused model is even better than Model 4 trained to maximize perceptual quality. It is worth noting that the bad convergence of Models 1 to 3 makes the traditional Adam/Adamw optimizer gain less improvement from the network fusion process. In contrast, the proposed hybrid optimization strategy facilitates the effectiveness of network fusion via the interaction of individuals in crossover operation. In addition, we find that AdamW does not show much advantages over Adam in the SR task. This is because AdamW is mainly proposed to handle the over-fitting problem in high-level vision tasks by introducing weight decay. However, this may not fit the SR task as the optimization might be under-fitting to the large space of image details.
@@
TABLE V: {justify} Quantitative comparison between the ‘Learnable Weight’, ‘Weight Regression’, and ‘Averaging Fusion’ on the Urban100 dataset. The RRDB backbone is adopted. The best and second results are highlighted in red and blue.
PSNR
SSIM
LPIPS
Learnable Weight
25.60
0.7748
0.1195
Weight Regression
25.71 0.7756 0.1124
Averaging Fusion
25.57
0.7752 0.1150
Network Fusion vs. Learnable Weight. We propose a two-stage framework for fusing $N$ networks. The first stage is the attention-based weight regression, which generates a set of weights for each input LR image. Therefore, the corresponding weight vectors of validation data can be obtained. The second stage is model fusion, which obtains a single stronger model by averaging all the weight vectors of validation data. To show the effectiveness of the proposed fusion method, we introduce a fusion method by setting the fusion weights of $N$ models as learnable parameters directly, which is named ‘Learnable Weight’. We compare our two stages, named ‘Weight Regression’ and ‘Averaging Fusion’, with ‘Learnable Weight’ on the Urban100 dataset in Table V. We can see that the ‘Weight Regression’ achieves the best results in all metrics since it can dynamically infer the optimal weights for each input based on image content. The ‘Averaging Fusion’ obtains better performance than ‘Learnable Weight’ because the universal network weight is inherited from ‘Averaging Fusion’ using validation data, which holds a stronger generalization capability.
TABLE VI: {justify} Quantitative comparison on Urban100 benckmark among different loss functions for the network fusion process. The RRDB backbone is adopted.
PSNR
SSIM
LPIPS
RRDB+Ours-$\ell_{1}$
26.55
0.8029
0.1955
RRDB +Ours-PD [65]
25.89
0.7824
0.1220
@@
Loss Selection for Network Fusion. The fusion network aims to leverage the strength of $N$ models for perceptually more plausible SR results. Table VI gives the numerical comparisons on the Urban100 benchmark with $\ell_{1}$, $PD$ [27] and $L_{adv}$ loss functions. We choose the adversarial loss in this paper because it achieves the best LPIPS score, which implies perceptually more realistic outputs.
@@
TABLE VII: {justify} Weights of two layers (L1 and L2) in the RRDB models obtained with different perception-distortion preferences given two inputs (I1 and I2), and the final fusion model (Ours).
L1
L2
I1
[0.0463, 0.0167, 0.0678, 0.3333, 0.5358]
[0.2348, 0.2084, 0.1426, 0.1452, 0.2689]
I2
[0.0263, 0.0144, 0.1014, 0.3307, 0.5272]
[0.3371, 0.1061, 0.0022, 0.1985, 0.3561]
Ours
[0.0965, 0.0733, 0.1429, 0.2629, 0.4243]
[0.3203, 0.1212, 0.0224, 0.1932, 0.3428]
@@
TABLE VIII: {justify} Quantitative comparisons on Urban100 benchmark between different ways of weight combination. The RRDB backbone is adopted. The best results are highlighted in bold.
PSNR
SSIM
LPIPS
21.86
0.6546
0.2269
24.52
0.7573
0.1755
25.57 0.7752 0.1150
@@
Weights of Obtained Models. To further show the advantage of our proposed method, we study on the weights of obtained models. Given two inputs (I1 and I2), the weights in two layers (L1 and L2) of the $5$ obtained RRDB models are shown in Table VII. The weights of the final fusion model ‘Ours’ are also given. It can be seen that the weights are neither evenly distributed among models nor concentrated on one or two models. The weight distribution varies in layers, and each model contributes to the final fusion. This demonstrates the effectiveness of our EA-Adam optimizer, which can simultaneously obtain models with different focuses. We also perform the ablation study on weight by (a) directly averaging the parameters of the $5$ networks, (b) applying the weights of one layer to all layers of the $5$ networks, and (c) our averaging fusion strategy. The results on Urban100 are shown in Table VIII. Our averaging strategy gets the best result, showing the effectiveness of our network fusion method.
@@
TABLE IX: {justify} Comparison of the training time of our hybrid EA-Adam optimization method and the Adam method.
@@
EA-iter
Adam-iter
Run-time/epoch
0.19h
4.76h
0.80h
0.63h
Total-epoch
250
8
92
50
Total-run-time
48h
112h
32h
IV-D Training and Inference Time
Training time. We first compare the training time of our hybrid EA-Adam optimization method (including the EA-Adam optimizer and the network fusion stage) and the original Adam optimizer by training an SR model with RRDB-backbone for ×4 super-resolution. Table IX provides the quantitative comparison of the training time. Specifically, the Adam optimizer costs 250 epochs and 48h to converge. Though hybrid EA-Adam costs a longer time in each epoch, it converges in much fewer epochs thanks to the effective interaction in the EA step. The hybrid EA-Adam optimization method totally costs 144h to finish the model training, which is about three times the training time of Adam. Most of the training time of hybrid EA-Adam is spent on the EA part. It is noted that there have been various parallel computation methods [73, 74] to speed up the EA process based on CUDA technology, showing great potential to accelerate our method in training.

### v2

@@ IV-C Ablation Studies
In this section, we conduct a series of ablation studies to validate the effectiveness of the proposed hybrid EA-Adam optimization. We first conduct experiments to compare our hybrid EA-Adam algorithm with the Adam algorithm, and compare their results with different selections of $N$. Then we compare our network fusion method with the learnable weight method. Furthermore, we study the selection of different loss functions in optimizing the weight regression network. We also study the aggregation weights of obtained models with different perception-distortion preferences. All ablation results are evaluated on the Urban100 dataset.
@@
Model Model1 Model2 Model3 Model4 Fused model
Weights $[0.25,0.75]$ $[0.5,0.5]$ $[0.75,0.25]$ $[1,0]$
PSNR Adam 24.36 23.82 23.96 24.22 24.41
EA-Adam 25.78 25.75 24.95 24.73 24.83
AdamW 25.48 24.67 23.41 24.18 24.37
EA-AdamW 25.95 25.81 25.21 24.82 24.91
LPIPS Adam 0.3061 0.2899 0.2665 0.1446 0.1437
EA-Adam 0.2120 0.2055 0.1536 0.1384 0.1354
AdamW 0.2570 0.2925 0.2306 0.1477 0.1444
EA-AdamW 0.2265 0.2201 0.1519 0.1404 0.1368
@@
Table IV gives the quantitative comparison. Without loss of generality, we conduct experiments with the SRResNet backbone and evaluate on the Urban100 dataset. PSNR and LPIPS metrics are used to measure the fidelity and perceptual quality, respectively. One can see that our hybrid EA-Adam/AdamW optimizers outperform Adam/AdamW by a large margin in most cases. Both the fidelity and perceptual quality metrics are improved, validating that the proposed hybrid EA-Adam strategy can push the PF closer to the optimal point of the metric space. For most models, EA-Adam improves PSNR by more than 0.4dB over Adam and improves LPIPS by more than $10\%$. This might be caused by the gradient vanishing problem in the gradient-based optimization process so that the models can be trapped into a local minimum. In contrast, with our hybrid EA-Adam optimizer, effective interaction between different models can be introduced by the crossover operator in EA, so that the models can be stably optimized with better convergence.
On the other hand, each of the four models may perform well in one aspect, yet sacrifices the other aspect as a trade-off. The fused model by our network fusion method achieves a better balance between fidelity and perceptual quality. Specifically, although the fidelity indices of the fused model are lower than that of Models 1, 2, and 3, its perceptual performance is better than Model 4, which is trained to maximize perceptual quality. This result well aligns with the goal of fusion network, which aims to further improve the perceptual quality of the fused model without reducing much the fidelity. It is worth noting that the bad convergence of Models 1 to 3 makes the traditional Adam/Adamw optimizer gain less improvement from the network fusion process. In contrast, the proposed hybrid optimization strategy facilitates the effectiveness of network fusion via the interaction of individuals in crossover operation. In addition, we find that AdamW does not show many advantages over Adam in the SR task. This is because AdamW is mainly proposed to handle the over-fitting problem in high-level vision tasks by introducing weight decay. However, this may not fit the SR task as the optimization might be under-fitting to the large space of image details.
@@
TABLE V: {justify} Training time and quantitative comparisons on Urban100 benchmark by using different numbers of individuals ($N$) in EA.
N PSNR SSIM LPIPS Training time
3 25.34 0.7525 0.1184 63h
5 25.57 0.7752 0.1150 112h
7 25.60 0.7804 0.1125 160h
9 25.64 0.7823 0.1114 211h
TABLE VI: {justify} Quantitative comparison between the ‘Learnable Weight’, ‘Weight Regression’, and ‘Our Fusion Method’ on the Urban100 dataset. The RRDB backbone is adopted. The best and second results are highlighted in red and blue.
PSNR SSIM LPIPS
Learnable Weight 25.60 0.7748 0.1195
Weight Regression 25.71 0.7756 0.1124
Our Fusion Method 25.57 0.7752 0.1150
Selection of the number of optimized models $N$. To show the influence of the number of optimized models, we conduct experiments by using different selections of $N$. The results on Urban100 are shown in Table V. One can see that increasing $N$ can obtain better perception-distortion performance but at the price of longer optimization time. When $N$ is greater than 5, the benefits are not significant. Considering both performance and training time, we choose $N=5$ in the experiments.
Network Fusion vs. Learnable Weight. We propose a two-stage framework for fusing $N$ networks. The first stage is the attention-based weight regression, which generates a set of weights for each input LR image. Therefore, the corresponding weight vectors of validation data can be obtained. The second stage is model fusion, which obtains a single stronger model by averaging all the weight vectors of validation data. To show the effectiveness of the proposed fusion method, we introduce a fusion method by setting the fusion weights of $N$ models as learnable parameters directly, which is named ‘Learnable Weight’. We compare our two stages, named ‘Weight Regression’ and ‘Our Fusion Method’, with ‘Learnable Weight’ on the Urban100 dataset in Table VI. We can see that the ‘Weight Regression’ achieves the best results in all metrics since it can dynamically infer the optimal weights for each input based on image content. ‘Our Fusion Method’ obtains better performance than ‘Learnable Weight’ in perceptual quality metric LPIPS with 0.0045 improvement while keeping comparable PSNR and SSIM values. This is because the universal network weight is inherited from ‘Weight Regression’ using validation data, which holds a stronger generalization capability.
TABLE VII: {justify} Quantitative comparison on Urban100 benckmark among different loss functions for the network fusion process. The RRDB backbone is adopted.
PSNR SSIM LPIPS
RRDB+Ours-$\ell_{1}$ 26.55 0.8029 0.1955
RRDB +Ours-PD [65] 25.89 0.7824 0.1220
@@
Loss Selection for Network Fusion. The fusion network aims to leverage the strength of $N$ models for perceptually more plausible SR results. Table VII gives the numerical comparisons on the Urban100 benchmark with $\ell_{1}$, $PD$ [28] and $L_{adv}$ loss functions. We choose the adversarial loss in this paper because it achieves the best LPIPS score, which implies perceptually more realistic outputs.
@@
TABLE VIII: {justify} Aggregation weights of two layers (L1 and L2) in the RRDB models with different perception-distortion preferences given two inputs (I1 and I2), and the final fusion model (Ours).
L1 L2
I1 [0.0463, 0.0167, 0.0678, 0.3333, 0.5358] [0.2348, 0.2084, 0.1426, 0.1452, 0.2689]
I2 [0.0263, 0.0144, 0.1014, 0.3307, 0.5272] [0.3371, 0.1061, 0.0022, 0.1985, 0.3561]
Ours [0.0965, 0.0733, 0.1429, 0.2629, 0.4243] [0.3203, 0.1212, 0.0224, 0.1932, 0.3428]
@@
TABLE IX: {justify} Quantitative comparisons on Urban100 benchmark between different ways of aggregation weight combination. The RRDB backbone is adopted. The best results are highlighted in bold.
PSNR SSIM LPIPS
(a) 21.86 0.6546 0.2269
(b) 24.52 0.7573 0.1755
(c) 25.57 0.7752 0.1150
@@
Aggregation Weights of Obtained Models. In the fusion stage, the network parameters of the $N$ models are frozen, and the aggregation weights are optimized for each layer using a set of validation data to fuse the $N$ models into one by interpolation, as done in those network interpolation methods [39]. To further show the advantage of our proposed fusion method, we study the aggregation weights of obtained models. Given two inputs (I1 and I2), the aggregation weights in two random layers (L1 and L2) of the $5$ obtained RRDB models are shown in Table VIII. The aggregation weights of the final fusion model ‘Ours’ are also given. It can be seen that the weights are neither evenly distributed among models nor concentrated on one or two models. The aggregation weight distribution varies in layers, and each model contributes to the final fusion. This demonstrates the effectiveness of our EA-Adam optimizer, which can simultaneously obtain models with different focuses. We also perform the ablation study on aggregation weight by (a) directly averaging the parameters of the $5$ networks, (b) applying the aggregation weights of one layer to all layers of the $5$ networks, and (c) our proposed fusion strategy. The results on Urban100 are shown in Table IX. Our fusion method achieves the best result, showing the effectiveness of our network fusion method.
IV-D Applications to real-world Super-Resolution
To demonstrate the generalization capability of the proposed method, we further present the comparisons by using the complex real-world SR setting. Specifically, we use the generator and discriminator of RealESRGAN [69] and optimize the model with our proposed EA-Adam hybrid optimizer and the fusion strategy. The DF2K [67, 68] is used as the training dataset, and the LR images are degraded with the RealESRGAN degradation pipeline [69]. The test data are from the real-world datasets, including RealSR [76] and DRealSR [77]. The compared methods include RealESRGAN [69], IKC [82], LDL [13], IKRNet [83] and DASR [38]. The results are shown in Table X. One can see that IKC obtains the worst results since it focuses on solving blur degradation and cannot adapt to real-world SR tasks. The proposed method obtains the best perception-distortion results in the real-world SR task, demonstrating its effectiveness.
TABLE X: {justify} Quantitative comparisons with GAN-based state-of-the-art real-world super-resolution methods on RealSR and DrealSR test datasets. The RRDB backbone is adopted. The best result is highlighted in bold.
RealSR DrealSR
PSNR/SSIM/LPIPS PSNR/SSIM/LPIPS
RealESRGAN 25.69/0.7616/0.2727 28.64/0.8053/0.2847
IKC 19.51/0.5088/0.4337 27.07/0.7429/0.3994
IKR-Net 25.15/0.7308/0.3912 28.94/0.8131/0.4274
LDL 25.28/0.7567/0.2766 28.21/0.8126/0.2815
DASR 27.02/0.7708/0.3151 29.77/0.8263/0.3126
RRDB+Ours 27.59/0.7806/0.2621 30.32/0.8358/0.2802
IV-E Training and Inference Time
Training time. We first compare the training time of our hybrid EA-Adam optimization method (including the EA-Adam optimizer and the network fusion stage) and the original Adam optimizer by training an SR model with RRDB-backbone for ×4 super-resolution. Table XI provides the quantitative comparison of the training time. Specifically, the Adam optimizer costs 250 epochs and 48h to converge. Though hybrid EA-Adam costs a longer time in each epoch, it converges in much fewer epochs thanks to the effective interaction in the EA step. The hybrid EA-Adam optimization method totally costs 144h to finish the model training, which is about three times the training time of Adam. Most of the training time of hybrid EA-Adam is spent on the EA part. It is noted that there have been various parallel computation methods [85, 86] to speed up the EA process based on CUDA technology, showing great potential to accelerate our method in training.
@@ IV-E Training and Inference Time
TABLE XI: {justify} Comparison of the training time of our hybrid EA-Adam optimization method and the Adam method.
@@
EA-iter Adam-iter
Run-time/epoch 0.19h 4.76h 0.80h 0.63h
Total-epoch 250 8 92 50
Total-run-time 48h 112h 32h
Figure 11: Visualization of failure cases generated by our model.
IV-F Failure Cases
While our method can improve the training process of GAN-based SR models and improve the model performance in perception-distortion balance, it can still fail in some challenging cases. Fig. 11 shows some typical examples. If the high-frequency structures are largely destroyed in the input LR image, it is hard for our model to correct them in the SR output, as shown in Fig. 11(a). In addition, for some small-scale textures and text characters, our method may generate some visual artifacts, as shown in Figs. 11(b)) and 11(c)).